OpenStudy (anonymous):

How do you complete the square with an equation like y = x^2 - 2x - 3? Please just show steps.

OpenStudy (anonymous):

All right, we can complete the square as such: \[ y=x^2-2x-3\\ \left(\frac{-2}{2}\right)^2=(-1)^2=1\implies\\ y=(x^2-2x+1)-3+1 \](We add unity outside since we just added unity inside) Factoring and simplifying gives us: \[ y=(x-2x+1)-2\implies\\ y=(x-1)^2-2 \]Et voilá.

OpenStudy (anonymous):

Dude thanks much. I was confused on the -2x because I thought you changed that. Thanks though.

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

Wait, my bad, I messed up on one thing, it is supposed to subtract one outside: \[ y=(x-1)^2-4 \]Not \(-2\).

OpenStudy (anonymous):

Yea I was just about to say. I'm trying to graph it right now and just noticed the vertex and the y-intercept doesn't match up.

OpenStudy (anonymous):

Yeah, here: \[ y=x^2-2x-3\\ \left(\frac{-2}{2}\right)^2=(-1)^2=1\implies\\ y=(x^2-2x+1)-3-1\implies\\ y=(x-1)^2-4 \]

OpenStudy (anonymous):

Couly you show me how you would do \[-3x ^{2} + 3x - 2\] please?

OpenStudy (anonymous):

Sure: \[ y=-3x^2+3x-2\implies\\ y=-3(x^2-x)-2 \]We know: \[ \left(-\frac{1}{2}\right)^2=\frac{1}{4} \]So: \[ y=-3\left(x^2-x+\frac{1}{4}\right)-2+3\cdot\frac{1}{4}=-3\left(x^2-x+\frac{1}{4}\right)-\frac{5}{4}\implies\\ y=-3\left(x-\frac{1}{2}\right)^2-\frac{5}{4} \]And, we're done.

OpenStudy (anonymous):

Thanks. It is starting to make since now.

OpenStudy (anonymous):

Cool, glad it is!